# Vulnerable Connectivity Caused by Local Communities in Spatial Networks

### v2

@@ Abstract
Local communities are widely observed in spatial networks. However, how such structure affects the vulnerability against malicious attacks remains unclear. This study investigates the impact of local communities on the robustness of connectivity by modeling planar infrastructure networks based on statistical population data. Our research reveals that the emergence of local communities is caused by spatial concentrations of nodes connected by short links, which significantly reduce the robustness. These results suggest that strategically establishing long-distance links provides a feasible solution to balance reliability and construction costs in infrastructure network design.
keywords:
Robustness of Connectivity, Local Community, Spatial Networks, Proximity Graphs, Population Distribution
@@ 1 Introduction
Community structure characterized by dense intra-community links and sparse inter-community bridge link is widely observed in various networks [1][2] including spatial networks [3][4][5][6]. Although there have been extensive researches on community detection [7][8], it still be not elucidated how community structure affects the vulnerability against attacks. We take notice to that such a structure creates potential vulnerabilities: When critical bridge links are removed, the network tends to fragment into components [9]. Previous research has shown that artificially adding community structure can significantly weaken the robustness of connectivity in scale-free (SF) networks and Erdős-Rényi (ER) random graphs [10][11]. Unlike only topological SF networks and ER graphs, spatial networks are constrained by physical and geographical factors, which may be related to community structure. As examples of these factors, in densely populated areas, road networks require more intersections to meet traffic demands[12], while wireless communication networks[13] need more base stations to maintain service quality. Although it has been pointed out that in a topological structure affected by geographical constraints on linking, the connectivity is weakened by constructing local stubs with small cycles [14], the reason for this weakness remains unclear without considering the influence of community structure.
@@
To address these gaps, we focus on road and communication networks as typical geographically constrained infrastructure networks. In constructing realistic spatial networks, we require reasonable data for locations of nodes. While there are many possible locations, the detail locations of nodes or links are not open usually for security reason to protect the using as potential targeted attacks. Therefore, we apply statistical population data of seven major urban areas in Japan as a realistic reference for determining node’s locations. Moreover, since spatial networks embedded on the surface of Earth have many short links and tend to be planar because of geographical limitations and economic considerations, we consider relative neighborhood graph (RNG) [15] and Gabriel graph (GG) [16] as modeling of road and communication networks, respectively. RNG and GG are studied in computer science, and have similarities to road networks [17] or the self-organized biological transport networks such as slime mould with interactions between the growing active zone and chemogradients [18].
@@
This paper aims to investigate how local communities affect the robustness of connectivity in spatial networks. In particular, we show that spatial concentrations of nodes connected by short links lead to the emergence of local communities, which make spatial networks more vulnerable against node removals.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce proximity graphs of RNG and GG with short links, which construct spatial networks based on population data that determines node locations. We consider three node removal strategies (1) recalculated betweenness attacks (RB) [19] by removing high betweenness centrality nodes, (2) initial degree attacks (ID) [20] by removing high degree nodes, and (3) random failure (RF) [9] by removing nodes randomly. We focus on RB, since betweenness centrality-based attacks are the most effective for fragmenting the largest connected component (LCC) in networks with communities, while ID and RF are typical node removals for comparing with the above damages. In Section 3, we show how local communities emerge and affect the robustness of connectivity. In particular, we conclude that spatial networks with strong community structure weaken the robustness against intentional attacks and random failures.
@@ 2.1 Models of Spatial Networks
Real networks such as road or communication systems tend to be planar on the surface of Earth. Thus, they are constrained by physical and geographic limitations. Economic considerations further reinforce this planarity for minimizing infrastructure investments. Moreover, the evolutionary nature of these networks by connecting neighbor nodes inherently leads to a planar structure through this proximity-based growth pattern. Based on this understanding, we introduce relative neighborhood graph (RNG) and Gabriel graph (GG) as models of road and communication networks, respectively, which are used to construct spatial networks. Note that RNG is a subgraph of GG and both of them are planar in proximity graphs with short links. These models have unique advantages for network analysis, particularly related to real-world population data as mentioned later.
RNG [15] and GG [16] are defined as follows. For a given set of nodes on a geographical space, two nodes are connected if no other node exists within a neighboring area of them. As shown in Figures 1 (a) (b), each shaded area is the intersection of two circles centered at two nodes with a radius equal to the distance between them.
RNG can be used to model road networks, because these links almost coincide with road segments in urban cities [21] [22]. GG has been used for modeling optical or ad-hoc wireless networks, because it is resemble to the structure of physical networks [16] [23] [24]. In particular, directional antennas are used in wireless networks with better beam power for linking [25]. In Figure 2, blue and orange shades represent strong area of directional beams, while the center circle represents the interference area. If other base station is located in this area with interference, connecting those two nodes is prohibited.
@@ 2.2 Node’s Locations and Removal Strategies
Taking into account the correlation between population density and node’s locations, we set three types of node’s locations: Population (Pop.), Inverse Population (Inv.) and Uniform (Uni.). Here, the location of nodes are determined by the centers of meshes according to statistical population data mentioned in the next section. For Pop., the nodes are located according to densely populated points ranked from highest to lowest, representing an extreme case where investments are exclusively concentrated in the most populated areas. For Inv., the nodes are located according to sparsely populated points ranked from the lowest to the highest, representing a strategic utilization of extremely sparse areas because of lower land prices for constructing networks with links by asphalt pavements or cables and node equipments such as traffic lights or communication devices. For Uni., the nodes are located according to a uniformly at random distribution, which moderate the concentration of nodes in Pop. and Inv. as these extreme cases. In most cities, population density is uneven, there is mixing of sparse and dense node’s locations in both Pop. and Inv.. The mixtures also occur in Uni. because of the spatial point process (SPP) theory [29]. Based on the SPP theory, the number of nodes in a given area follows a Poisson distribution, therefore there are mixtures of sparse and dense node’s locations. These mixtures give local community structure as shown later. By connecting nodes on the locations of Pop., Inv., and Uni., spatial networks are constructed according to the previously described constraints of short links in RNG and GG shown in Figure 1. Moreover, as completely uniform without mixtures of dense and sparse node’s locations, we consider the fourth type of node’s locations on a 2D square lattice (2DL) [30] in matching the condition under a same $P(k)$ for comparing the robustness in 2DL and Pop., Inv., or Uni. networks. According to the slightly different degree distributions $P(k)$ as shown later obtained from Pop., Inv., and Uni. networks, we randomly assign the corresponding number of links to each node on the 2DL. We assume that the other end of assigned link is free (unconnected) to perform the following two steps of trials for linking. The first trial is established by connecting free-ends between neighbor nodes on the lattice. When free-ends are remaining but its neighbor nodes have no available free-ends, the second trial is performed for connecting more distant nodes with free-ends available. Note that the 2DL is not planar with the second trial links.
@@
On the other hand, as node removals, we consider three typical strategies: recalculated betweenness (RB) attacks [19], initial degree (ID) [20] attacks, and random failures (RF) [9]. In these strategies, nodes (and links connected to them) are removed one by one until $q/N$ nodes are removed, where $q$ and $N$ denote the fraction of removed nodes and network size, respectively. Removed nodes are selected in decreasing order of degrees for ID attacks or betweenness centrality for RB attacks. Note that RB attacks recalculate betweenness centrality, while ID attacks select nodes according to the initial degree. Besides, RF is considered to enhance the difference between intentional attacks and random failures. We emphasize that RB has a strong affect on global fragmentation in networks with communities.
@@ 3 Effects of Community Structure on the Robustness
This section numerically shows important results for the robustness in the spatial networks. In subsection 3.1, we show the emergence of local communities in RNG and GG as modeling road and communication networks whose nodes are located by Pop., Inv., and Uni.. In subsections 3.2 and 3.3, we show that spatial concentrations of nodes connected by short links make local communities, which affect the robustness of connectivity against both intentional attacks and random failures.
@@ 3.1 Emergence of Local Communities
As real population data, the seven Japanese areas used in our study are Fukuoka, Hiroshima, Kyoto-Osaka (referred to as Keihan), Nagoya, Tokyo, Sendai, and Sapporo [31]. Each of them is divided into $500m$ $500m$ block meshes with totaling (8(8( 8 × $10$ $2$ $3)^{2}=230,400$ blocks, while nodes are located at the centers of meshes. We select $N$ nodes whose locations are based on Pop., Inv. and Uni., respectively, as mentioned in subsection 2.2 previously. To ensure that $\sqrt{N}$ is an integer when constructing a 2DL for later comparison, we set $N=100,1024$ and $10000$. Links between nodes are created by connection rules of RNG and GG as shown in Figure 1. Local communities emerge in such spatial networks.
To measure the strength of community structure, we introduce modularity $Q$ [32] defined as:
@@
where $A_{ij}$ denotes the $i$, $j$ element of adjacency matrix: $A_{ij}=1$ if there is a link between nodes $i$ and $j$, 0 otherwise. $k_{i}$ and $k_{j}$ are the degrees of them, respectively. $M$ is the total number of links, $c_{i}$ represents the community identifier to which node $i$ belongs, $\delta(c_{i},c_{j})$ is an indicator function equal to 1 if nodes $i$ and $j$ belong to a same community, and 0 otherwise. We estimate $c_{i}$ by using Louvain method [33]. A higher value of $Q$ indicates a stronger community structure, where there are many intra-links however a few inter-links.
@@
We explain the typical results of communities for Tokyo $N=1024$ node networks as examples. For other cities, the results are shown in Supporting Information from S1 Figure to S12 Figure. Figures 3 and 4 show the communities for node’s location based on Pop. and Inv. with short links created by RNG and GG. Since RNG is a subgraph of GG, the numbers of links in Figures 3 and 4 (a) are less than them in Figures 3 and 4 (b). Thus, the results of communities in RNG and GG networks are different. Moreover, Figures 5 and 6 show the examples of communities for node’s location based on Pop. and Inv. after 10% nodes (and links connected to them) removals against RB attacks. The removal of nodes causes the network to break into disconnected components, particularly at the boundaries between different communities. This fragmentation is the most noticeable in densely populated areas. For all seven Japanese areas, Table 1 shows that both Pop. and Inv. have slightly higher $Q$ values compared to Uni. for $N=1024$. This trend is also observed for other sizes $N$ in Supporting Information S6 Table and S7 Table. These results suggest that Pop. and Inv. networks have slightly stronger community structure than that of Uni. networks according to these spatial concentrations of nodes. Although Uni. has a slightly lower $Q$ value than that of Pop. and Inv., it still maintains community structure obviously. Remember that from the SPP theory in subsection 2.2, there is a mixture of sparse and dense node’s locations, and contributes to the emergence of community structure even in Uni. networks.

### v4

@@ Abstract
Local communities by concentration of nodes connected with short links are widely observed in spatial networks. However, how such structure affects robustness of connectivity against malicious attacks remains unclear. This study investigates the impact of local communities on the robustness by modeling planar infrastructure networks whose node’s locations are based on statistical population data. Our research reveals that the robustness is weakened by strong local communities in spatial networks. These results highlight the potential of long-distance links in mitigating the negative effects of local community on the robustness.
@@ 1 Introduction
Community structure characterized by dense intra-community links and sparse inter-community bridge link is widely observed in various networks [1][2], including spatial networks [3][4][5][6]. There are huge methods to detect communities as summarized in survey papers [7][8]. One of the most popular approaches is modularity maximization, while it has the limitations for hierarchical structure or overlapping among communities [9]. Thus, other approaches have been proposed such as statistical inference framework that combines the stochastic block model with belief propagation [10] [11] [12], spin-glass models that minimize an energy [13], and spectral methods [14] based on the non-backtracking matrix for sparse networks [15]. They are applied to real-world networks including chromatin maps [16], neural connectomes [17], and cryptocurrency markets, where spectral modularity also detects core–periphery structures [18]. Such extensive researches exist for community detection, however, it has not yet been elucidated how community structure affects the robustness of connectivity against attacks. From only topological point of view, artificially adding community structure is significantly weaken the robustness of connectivity in scale-free (SF) networks and Erdős-Rényi (ER) random graphs [19][20]. Although it has been pointed out that in a topological structure affected by geographical constraints on linking, the connectivity is weakened by constructing local stubs with small cycles [21], the reason for this weakness remains unclear without considering the influence of community structure.
@@
Unlike only topological SF networks and ER graphs, spatial networks are constrained by physical and geographical factors, which give locally concentration of nodes and relatively short links between nodes. As examples of these factors, in densely populated areas, road networks require more intersections to meet traffic demands [22], while wireless communication networks [23] need more base stations to maintain service quality. However, the detail locations of nodes or links are not open usually for security reason to protect the using as potential targeted attacks. Therefore, we simplify road and communication networks in proper abstraction. Since spatial networks embedded on the surface of Earth have many short links and tend to be planar [24], we use relative neighborhood graph (RNG) [25] and Gabriel graph (GG) [26] as models for road and communication networks, respectively. RNG and GG are known as planar proximity graphs with short links in computer science, and have similarities to road networks [27] or the self-organized biological transport networks such as slime mould with interactions between the growing active zone and chemogradients [28]. On the other hand, node’s locations should be set as real as possible. Thus, we apply statistical population data to determine node’s locations.
While robustness usually refers to the ability of a network to maintain connectivity under failures, resilience is a broader concept increasingly recognized as vital for infrastructure systems. It involves not only the capacity to resist and absorb disruptions, but also to recover and adapt afterward [29] [30]. In spatial networks, resilience has been modeled using cascading failure processes [31], percolation thresholds [32], and more recently, spatial association metrics that capture inter-regional recovery dynamics [33]. Though rich in implications, resilience-oriented approaches often rely on assumptions regarding functional loads, dependency links, or time-evolving recovery speeds. In this study, we limit our scope to static structural robustness as robustness of connectivity shaped by spatial node’s locations. However, the study of resilience in our networks will be a promising direction for future work.
@@
This paper aims to investigate how local communities (will be explained later) affect the robustness of connectivity in spatial networks. In particular, we show that spatial concentrations of nodes connected by short links lead to the emergence of local communities, which make spatial networks more vulnerable against node removals.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce proximity graphs of RNG and GG with short links, which construct spatial networks whose node’s locations are based on population data. We consider three node removal strategies (1) recalculated betweenness attacks (RB) [34] by removing high betweenness centrality nodes, (2) initial degree attacks (ID) [35] by removing high degree nodes, and (3) random failure (RF) [36] by removing nodes randomly. We focus on RB, since betweenness centrality-based attacks are the most effective for fragmenting the largest connected component (LCC) in networks with communities, while ID and RF are typical node removals for comparing with the above damages. In Section 3, we show how local communities emerge and affect the robustness of connectivity. In particular, we conclude that spatial networks with strong community structure weaken the robustness against intentional attacks and random failures.
@@ 2.1 Models of Spatial Networks
Most spatial networks such as infrastructure networks can be approximated by planar networks [37]. For example, drivable and walkable street networks in 50 global cities are investigated by using two planarity metrics of spatial planarity ratio $\phi$, which measures the proportion of true intersections to apparent ones, and the edge length ratio $\lambda$, which compares the mean edge lengths in planar and non-planar representations [24]. The study finds that cities in Europe (e.g., Florence, Paris) and China (e.g., Xi’an, Guangzhou) have higher planarity, while some U.S. cities such as Los Angeles and Dallas are less planar. However, most cases of $\phi$ and $\lambda$ are larger than 0.80. The results suggests that planar models give good approximating for urban street networks. A similar assumption can be extended to parts of backbone communication networks, because the fiber-optic cables are usually embedded under the main trunk roads.
Although real infrastructure systems may consist of multiple interdependent layers, such as power-grids and communication networks [38], [39], we restrict our modeling to single-layer spatial networks. This simplification enables us to isolate the effect of community structures formed by spatial node’s locations on the robustness of connectivity. Incorporating multilayer interactions would require assumptions about inter-layer dependencies, which fall beyond the scope of this structural analysis and are left for future exploration.
It has been shown that RNG can effectively approximate the spatial structure of urban road networks in both Japanese and American cities [40] [41]. This planarity results from a combination of physical constraints, cost-efficiency considerations, and the growth processes by connecting neighbor nodes in spatial networks. Since spatial networks can be approximated by planar networks, we introduce RNG and GG as models of road and communication networks, respectively, Note that RNG is a subgraph of GG and both of them are planar in proximity graphs with short links. Here, GG has been used for modeling optical or ad-hoc wireless networks, because it resembles to the structure of physical networks [26] [42] [43] partially. While many proximity-based or planar graphs such as Delaunay triangulations [44], k-nearest neighbor (k-NN) graphs [25], Voronoi-based constructions [45] and others exist, we focus on the Gabriel Graph (GG) and its subgraph of Relative Neighborhood Graph (RNG). We emphasize that the obtained results will be applied to a wide class of planar proximity networks, which include power-grids, water supply, gas pipline, and railway networks on the surface of Earth. Moreover, it has been shown that RNG effectively approximates real-world urban street networks [40, 41].
Fig. 1 RNG and GG. Illustration of connection constraints for (a) RNG and (b) GG. A link colored by red is established between two nodes colored by blue, when no other node exists within the shaded area.
RNG [25] and GG [26] are defined as follows. For a given set of nodes on a geographical space, two nodes are connected if no other node exists within a neighboring area of them. As shown in Figs. 1 (a) (b), each shaded area is the intersection of two circles centered at two nodes with a radius equal to the distance between them. In Fig. 2, blue and orange shades represent strong area of directional beams, while the center circle represents the interference area. If other base station is located in this area with interference by directional beams [46], connecting those two nodes is prohibited. All pairwise distances are computed using Euclidean metrics.
Fig. 2 Signal Coverage. Coverage diagram of radio waves between two base stations in wireless communication. The ranges of strong beams are shown by blue and orange shades. The center circle represents the signal interference area. If other base stations exist within it, the two stations cannot be connected.
@@ 2.2 Node’s Locations and Removal Strategies
Taking into account the correlation between population density and node’s locations, we set three types of node’s locations: Population (Pop.), Inverse Population (Inv.) and Uniform (Uni.). In a $480\times 480$ grid of $500\mathrm{m}\times 500\mathrm{m}$ block meshes, each mesh has residential population as mentioned later. For Pop. and Inv., all zero-population blocks are excluded. Here, a node is located by the center of mesh. For Pop., the nodes are located according to densely populated points ranked from highest to lowest, representing an extreme case where investments are exclusively concentrated in the most populated areas. For Inv., the nodes are located according to sparsely populated points ranked from the lowest to the highest, representing a strategic utilization of extremely sparse areas because of lower land prices for constructing networks with links by asphalt pavements or cables and node equipments such as traffic lights or communication devices. This ranking-based node selection is motivated by the observed population distribution of real cities. As shown in Supporting Information S38 Fig, when block meshes are ranked in decreasing order by population, the tails of distributions follow exponential decay, where only a small fraction of meshes contain most of the total population. Therefore, selecting the top or bottom $N$ meshes allows us to represent two statistically extreme cases of strong concentration versus wide dispersion without introducing artificial thresholds. For Uni., the nodes are located according to a uniformly at random distribution, which moderate the concentration of nodes in Pop. and Inv. as these extreme cases. In most cities, population density is uneven, there is mixing of sparse and dense node’s locations in both Pop. and Inv.. The mixtures also occur in Uni. because of the spatial point process (SPP) theory [50]. Based on the SPP theory, the number of nodes in a given area follows a Poisson distribution, therefore there are mixtures of sparse and dense node’s locations. These mixtures give local community structure as shown later. By connecting nodes on the locations of Pop., Inv., and Uni., spatial networks are constructed according to the previously described constraints of short links in RNG and GG shown in Fig. 1. Moreover, as completely uniform without mixtures of dense and sparse node’s locations, we consider the fourth type of node’s locations on a 2D square lattice (2DL) [51] in matching the condition under a same $P(k)$ for comparing the robustness in 2DL and Pop., Inv., or Uni. networks. According to the slightly different degree distributions $P(k)$ as shown later obtained from Pop., Inv., and Uni. networks, we randomly assign the corresponding number of links to each node on the 2DL. We assume that the other end of assigned link is free (unconnected) to perform the following two trials for linking. The first trial is established by connecting free-ends between neighbor nodes on the lattice. When free-ends are remaining but its neighbor nodes have no available free-ends, the second trial is performed for connecting more distant nodes with free-ends available. Note that the 2DL is not planar with the second trial links.
On the other hand, in practical scenarios, infrastructure networks may be exposed to various types of disruptions, such as natural disasters (e.g., earthquakes, floods), malicious attacks (e.g., cyber-attacks on communication systems, deliberate closure of key intersections), and operational failures (e.g., congestion at major road junctions or overloads in central routers). To investigate the impact of such disruptions on network connectivity, we consider three typical node removals that tend to be chosen by terrorists or to be exposed in disasters. Recalculated betweenness (RB) attacks [34] simulate targeted disruptions of nodes that carry high levels of networks flows, like busy intersections in road systems or central equipments in communication infrastructures. Here, for a node $v$, the betweenness centrality $b(v)$ [52] is defined as:
$b(v)=\sum_{s\neq v\neq t}\frac{\sigma_{st}(v)}{\sigma_{st}},$
@@
where $\sigma_{st}$ is the total number of shortest paths between nodes $s$ and $t$, and $\sigma_{st}(v)$ is the number of those paths that pass through $v$. Initial degree (ID) attacks [35] represent failures at nodes with a large number of direct connections, which are more likely to become congested because of their high visibility, like urban traffic hubs or widely connected access routers. Random failure (RF) attacks [36] model unintentional and spatially uncorrelated disruptions, such as equipment malfunctions, local construction damage, or accidental shutdowns that may occur anywhere in the network without targeting specific node properties. In these strategies, nodes (and links connected to them) are removed one by one until $q/N$ nodes are removed, where $q$ and $N$ denote the fraction of removed nodes and network size, respectively. Removed nodes are selected in decreasing order of degrees for ID attacks or betweenness centrality for RB attacks. Note that in RB attacks, betweenness centrality is recalculated after each node removal, while ID attacks select nodes according to the initial degree. Besides, RF is considered to enhance the difference between intentional attacks and random failures. We emphasize that RB has a strong affect on global fragmentation in networks with communities.
Other attacks, such as cascading failures and localized attacks (LA), are not considered in this study. Cascading failures simulate overload redistribution and recursive node collapses, commonly used to model flow-based infrastructure disruptions [31] [53] [54]. Localized attacks by target nodes within a confined spatial region and may induce system-wide collapse when the affected area exceeds a critical radius. It is shown that LA on spatially embedded networks with dependencies can trigger cascading-like propagation, leading to abrupt global fragmentation even when the initial damage is spatially limited [55]. While these approaches provide valuable insights into failure dynamics, they require further assumptions about flow, thresholds, or spatial radius. Given our focus on structural vulnerability shaped by node placement and short-range connectivity, we restrict our analysis to static removal strategies and leave these dynamic or spatially constrained models for future work.
@@ 3 Effects of Community Structure on the Robustness
This section numerically shows important results for the robustness in the spatial networks. In subsection 3.1, we show the emergence of local communities in RNG and GG as modeling road and communication networks whose node’s locations are based on Pop., Inv., and Uni.. In subsections 3.2 and 3.3, we show that spatial concentrations of nodes connected by short links make local communities, which affect the robustness of connectivity against both intentional attacks and random failures.
@@ 3.1 Emergence of Local Communities
As real population data, the seven Japanese areas used in our study are Fukuoka, Hiroshima, Kyoto-Osaka (referred to as Keihan), Nagoya, Tokyo, Sendai, and Sapporo [56]. Each of them is divided into $500m$ $500m$ block meshes with totaling (8(8( 8 × $10$ $2$ $3)^{2}=230,400$ blocks. We adopt $500\mathrm{m}\times 500\mathrm{m}$ meshes because this is the finest spatial resolution available in dataset delivered from the Japanese census. Nodes are located at the centers of meshes. We select $N$ nodes whose locations are based on Pop., Inv. and Uni., respectively, as mentioned in subsection 2.2 previously. To ensure that $\sqrt{N}$ is an integer when constructing a 2DL for later comparison, we set $N=100,1024$ and $10000$. Links between nodes are created by connection rules of RNG and GG as shown in Fig. 1. Local communities emerge in such spatial networks. In this paper, we refer to local communities as topological modules estimated by community detection methods. For example, Fig. 3a shows the detected community of Tokyo, whose links are constructed using RNG. The blue-colored community covers the area around Chiyoda and Chūō wards with the imperial palace and central government institutions as forming the central administrative and political core in Japan. The green-colored community extends southward into Minato and Shinagawa wards known for their high-end commercial zones and international transit hubs including Shimbashi and Shinagawa stations in Tokyo city. This pattern shows that local communities estimated in proximity-based spatial networks reflect meaningful urban divisions, which may be not shaped exactly by administrative orders but by population distribution and geographic connectivity. To measure the strength of community structure, we introduce modularity $Q$ [57] defined as:
@@
where $A_{ij}$ denotes the $i$, $j$ element of adjacency matrix: $A_{ij}=1$ if there is a link between nodes $i$ and $j$, 0 otherwise. $k_{i}$ and $k_{j}$ are the degrees of them, respectively. $M$ is the total number of links, $c_{i}$ represents the community identifier to which node $i$ belongs, $\delta(c_{i},c_{j})$ is an indicator function equal to 1 if nodes $i$ and $j$ belong to a same community, and 0 otherwise. We estimate $c_{i}$ by using the popular Louvain method [58] [59] [18], which maximizes the modularity $Q$ using only node degrees as local information. While modularity $Q$ provides a well-defined approach to measure the strength of community, it is not suitable for detecting overlapping or hierarchical communities [11], and its maximization is known to be NP-hard [60]. However, our goal is not to achieve precise community detection, but to compare the relative strength of community in different spatial networks. As a complementary measure, we also investigate the sparsity of networks, which will be detailed in the following subsection.
@@
We explain the typical results of estimated communities for Tokyo $N=1024$ node networks by Louvain method. For other cities, the results are shown in Supporting Information S1 Fig to S12 Fig. Figs. 3ab and 4ab show the communities of networks constructed by using node’s locations based on Pop. and Inv. with short links constrained by RNG and GG. Moreover, Figs. 3cd and 4cd show the communities of networks by using node’s locations based on Pop. and Inv. after 10% nodes (and links connected to them) removals against RB attacks. The removal of nodes causes the network to break into disconnected components, particularly at the boundaries between different communities. This fragmentation is the most noticeable in densely populated areas. For all seven Japanese areas, Table 1 shows that, in the original networks, both Pop.-based and Inv.-based networks have slightly higher $Q$ values compared to Uni.-based networks. Similarly, in the 2D lattice, most cases of Pop.-based and Inv.-based networks have slightly higher $Q$ values compared to Uni.-based networks, while exceptions exist in Pop-based networks that links are connected with RNG, where $Q$ values are slightly lower than those of Uni.-based networks (also see community number in S11 Table). A similar ordering of $Q$ values, where Pop.-based and Inv.-based networks higher than Uni.-based networks, is also observed for other networks sizes $N$ as shown in Supporting Information S6 Table and S7 Table. Differences in $Q$ of Pop.-based, Inv.-based, and Uni.-based networks suggest that Pop.-based and Inv.-based networks usually have slightly stronger local community than Uni.-based networks, which resulting from the spatial concentrations of nodes. Although Uni.-based networks have slightly lower $Q$ values than that of Pop.-based and Inv.-based networks, they still maintain community structure obviously. Remember that from the SPP theory in subsection 2.2, there is a mixture of sparse and dense node’s locations, and contributes to the emergence of community structure even in Uni.-based networks. Note that modularity $Q$ of Pop.-based networks relocated on a 2D lattice is smaller than that of Inv.-based and Uni.-based networks in RNG. While spatial density of node’s locations in Pop.-based, Inv.-based, and Uni.-based networks are weaken in the corresponding 2D lattice networks under the same degree distributions, respectively. We will explain how degree distribution affects the robustness of connectivity later.
Fig. 3 Community Structure of Tokyo Pop.-based Network. Visualization of community structure in (left) RNG and (right) GG (a)(b) as the original and (c)(d) after 10% nodes removal for Tokyo with $N=1024$ nodes located decreasing order of population (Pop.). Note that RNG is a subgraph of GG. Different communities detected by the Louvain method are colored by red, yellow, green, and blue. Each figure depicts a spatial network on a two-dimensional Euclidean plane, where node’s locations correspond to the centroid locations of 500m × 500m population mesh blocks. The horizontal and vertical axes represent planar distances derived from these centroid positions.
Fig. 4 Community Structure of Tokyo Inv.-based Network. Visualization of community structure in (left) RNG and (right) GG (a)(b)as the original and (c)(d) after 10% nodes removal for Tokyo with $N=1024$ nodes located inverse decreasing order of population (Inv.). Note that RNG is a subgraph of GG. Different communities detected by the Louvain method are colored by red, yellow, green, and blue. Each figure depicts a spatial network on a two-dimensional Euclidean plane, where node’s locations correspond to the centroid locations of 500m × 500m population mesh blocks. The horizontal and vertical axes represent planar distances derived from these centroid positions.
Table 1: Modularity $Q$ in networks with 1024 nodes for seven major Japanese areas under the original and 2D lattice networks
Cities Original 2D lattice
RNG GG RNG GG
Inv. Pop. Inv. Pop. Inv. Pop. Inv. Pop.
Fukuoka 0.8905 0.8953 0.8728 0.8857 0.8726 0.8298 0.7677 0.8269
Hiroshima 0.8866 0.8993 0.8717 0.8898 0.8704 0.8332 0.7662 0.8203
Keihan 0.8903 0.8924 0.8651 0.8777 0.8693 0.8302 0.7756 0.8149
Nagoya 0.8880 0.8929 0.8612 0.8764 0.8807 0.8408 0.7744 0.8158
Tokyo 0.8846 0.8814 0.8615 0.8610 0.8644 0.8301 0.7548 0.8177
Sendai 0.8853 0.8908 0.8612 0.8801 0.8653 0.8387 0.7648 0.8184
Sapporo 0.9017 0.8775 0.8875 0.8722 0.8745 0.8187 0.7863 0.8144
Uniform 0.8735 0.8557 0.8649 0.7450
Modularity $Q$ in networks before (original) and after node relocation on a 2D lattice (relocated). Higher values indicate stronger community structure. Values with upper-triangles ($\triangle$) indicate that Pop.-based or Inv.-based networks have higher modularity than Uni.-based networks in both RNG and GG. Bold values indicate networks have lower modularity than that of Uni.-based networks.